\bigskip

\begin{thebibliography}{}

\bibitem{mg79b} E.P. Mazets, S.V. Golenetskii, V.N. Il’Inskii, V.N. Panov, R.L. Aptekar’, et al (1979) A flaring x-ray
pulsar in Dorado. Sov Astron Lett 5:163–166

\bibitem{mg79c} E.P. Mazets, S.V. Golenetskii, V.N. Il’Inskii, R.L. Aptekar’ Yu.A. Guryan (1979) Observations of a flaring
X-ray pulsar in Dorado. Nature 282:587–589

\bibitem{gm79}  S.V. Golenetskii, E.P. Mazets, V.N. Il’Inskii, Yu.A. Guryan (1979) Recurrent gamma ray bursts from the
flaring source FXP 0520 - 66. Sov Astron Lett 5:340–342

\bibitem{mg82}  E.P. Mazets, S.V. Golenetskii, Yu.A. Guryan, V.N. Il’Inskii (1982) The 5 March 1979 event and the distinct class of short gamma bursts: are they of the same origin. Astrophys Space Sci 84:173–189

\bibitem{ruf17} R.C.R. de Lima, J.G. Coelho, M. Malheiro,
J.A. Rueda, and R. Ruffni (2017) SGRs/AXPs as Rotation-Powered Neutron Stars. IJMP: Conference Series Vol. 45 (2017) 1760030 (7 pages)

\bibitem{bk99} G.S. Bisnovatyi-Kogan (1999) Magnetic fields of neutron stars: very low and very high. Proceedings of the Vulcano Workshop 1999, Memorie della Societa Astronomica
Italiana 73:318–329 (2002);(arXiv:astroph/ 9911275)

\bibitem{ma08} E.P. Mazets, R.L. Aptekar’, T.L. Cline, D.D. Frederiks, J.O. Goldsten, et al (2008) A Giant Flare from a Soft Gamma Repeater in the Andromeda Galaxy (M31). ApJ 680:545–549

\bibitem{fp07b} D.D. Frederiks, V.D. Palshin,  R.L. Aptekar’, S.V. Golenetskii, T.L. Cline, et al (2007) On the possibility
of identifying the short hard burst GRB 051103 with a giant flare from a soft gamma repeater
in the M81 group of galaxies. Astron Lett 33:19–24

\bibitem{plp05}
A. Pozanenko, V. Loznikov, R. Preece (2005) Possible connection of soft gamma-repeaters  with short duration GRB. In: Proceedings of the XLth Rencontres de Moriond, eds.J. Dumarchez and Jean Tran Thanh, 253

\bibitem{hbs05}  K. Hurley, S.E. Boggs, D.M. Smith, et al (2005)  An exceptionally bright flare from SGR 1806-20 and the origins of short-duration $\gamma$-ray bursts. Nature 434:1098-1103

\bibitem{mc99c} E.P. Mazets, T.L. Cline, R.L. Aptekar’, 
 P. Butterworth, D.D. Frederiks, et al (1999) Activity of the soft gamma repeater SGR 1900+14 in 1998 from KONUS-Wind  observations: 2. The giant August 27 outburst. Astron Lett 25:635–648

\bibitem{mc05}  E.P. Mazets, T.L. Cline, R.L. Aptekar’, D.D. Frederiks, S.V. Golenetskii, et al (2005) The KONUS-Wind
and Helicon-Coronas-F detection of the giant gamma ray flare from the soft gamma ray repeater
SGR 1806-20. arXiv:astro-ph/0502541

\bibitem{ma99a} E.P. Mazets, R.L. Aptekar’, P. Butterworth,  T.L. Cline, D.D. Frederiks,
 et al (1999) Unusual Burst
Emission from the New Soft Gamma Repeater SGR 1627-41. ApJ Lett 519:L151–L153

\bibitem{dt92}  R.C. Duncan, C. Thompson (1992) Formation of very strongly magnetized neutron stars –implications for gamma ray bursts. ApJ Lett 392:L9–L13

\bibitem{dt95} R.C. Duncan, C. Thompson (1995) The soft gamma repeaters as very strongly magnetized neutron stars – I. Radiative mechanism for outbursts. MNRAS 275:255–300

\bibitem{bkch74} G.S. Bisnovatyi-Kogan, V.M. Chechetkin (1974) Nucleosynthesis in supernova outbursts and the chemical composition of the envelopes of neutron stars. ApSS 26:25–46

\bibitem{bkinch75} G.S. Bisnovatyi-Kogan, V.S. Imshennik, D.K. Nadyozhin, V.M. Chechetkin (1975) Pulsed gamma ray emission from neutron and collapsing stars and supernovae.
 ApSS 35:23–41

\bibitem{bkch81} G.S. Bisnovatyi-Kogan, V.M. Chechetkin (1981) Gamma ray bursts as a manifestation of the inherent activity of neutron stars. Sov Astron 58:561–568

\bibitem{bk92} G.S. Bisnovatyi-Kogan (1992) Gamma ray bursts from nuclear fission in neutron stars. In: Proceedings of the conference “Gamma Ray bursts – observations, analyses and theories”, July
29-August 3, 1990, Taos, New Mexico, USA. Cambridge University Press, pp 89–98

\bibitem{bk12} G.S. Bisnovatyi-Kogan (2012) Gamma ray bursts, their afterglows, and soft gamma repeaters. In: Proceedings of the workshop gamma ray bursts: probing the science, progenitors and their environment, Moscow, 13–5 June 2012, pp 1–4. 
    http://www.exul.ru/workshop2012/Proceedings2012.pdf

\bibitem{ks99} C. Kouveliotou, T. Strohmayer, K. Hurley, J. van Paradijs, M.H. Finger, et al (1999) Discovery of a Magnetar Associated with the Soft Gamma Repeater SGR 1900+14. ApJ Lett 510:L115–L118

\bibitem{mg79a}  E.P. Mazets, S.V. Golenetskii, Yu.A. Guryan,
 (1979) Soft gamma ray bursts from the source B1900+14. Sov Astron Lett 5:343–344

\bibitem{mc99b}  E.P. Mazets, T.L. Cline, R.L. Aptekar’, P. Butterworth D.D. Frederiks, et al (1999) Activity of the soft gamma repeater SGR 1900+14 in 1998 from KONUS-Wind observations: 1. Short recurrent bursts. Astron Lett 25:628–634

\bibitem{pb05} D.M. Palmer, S. Barthelmy, N. Gehrels, R.M. Kippen, T. Cayton, et al (2005) A giant gamma ray flare from the magnetar SGR 1806-20. Nature 434:1107–1109

\bibitem{kd98}  C. Kouveliotou, S. Dieters, T. Strohmayer, J. van Paradijs, G.J. Fishman, et al (1998) An X-ray pulsar with a superstrong magnetic field in the soft gamma ray repeater SGR 1806-20. Nature 393:235–237

\bibitem{uf93} A. Ulmer, E.E. Fenimore, R.I. Epstein, C. Ho, R.W. Klebesadel, J.G. Laros, F. Delgado (1993) Soft Gamma Repeater 1806-20: Weak Evidence for a Periodicity following the Bursts.  ApJ 418:395-397

\bibitem{lf86a} J.G. Laros, E.E. Fenimore, R.W. Klebesadel, S.R. Kane (1986) Statistical Properties of the Soft Gamma Repeater (SGR) GB790107. Bull Am Astron Soc 18:928

\bibitem{lf86b} J.G. Laros, E.E. Fenimore, M.M. Fikani, R.W. Klebesadel, C. Barat (1986) The soft gamma ray burst GB790107. Nature 322:152–153

\bibitem{re10} N. Rea, Esposito P., Turolla R., Israel G.L., Zane S. et al (2010) A low-magnetic-field soft gamma repeater. Science 330:944–946

\bibitem{aec11} M.A. Alpar, U. Ertan, and S. Calıskan (2011) SGR 0418+5729 — How does a young neutron star spin down to a 9 s period with a dipole field less than 1013 G? Astrophysical Journal Letters 732:L4 (4pp)

\bibitem{gbk20} E. Göğüş, M.G. Baring, C. Kouveliotou, et al (2020)
Persistent Emission Properties of SGR J1935+2154 during its 2020 active episode. Astrophysical Journal Letters 905:L31 (6pp)

\bibitem{msk03}  M.A. McLaughlin, I.H. Stairs, V.M. Kaspi, D.R. Lorimer, M. Kramer, et al (2003) PSR J1847–0130: A radio pulsar with magnetar spin characteristics. ApJ Lett 59l:L135–L138

\bibitem{bk16} G.S. Bisnovatyi-Kogan (2016) Soft gamma repeaters/anomalous X-ray pulsars – are they magnetars? Astron Astrophys Trans 29:165–182

\bibitem{bk17} G.S. Bisnovatyi-Kogan (2017) Young neutron stars with soft gamma ray emission, and anomalous X-Ray pulsars. Handbook of Supernovae.  
    Springer International Publishing AG,  pp 1401-1423

\bibitem{bki14} G.S. Bisnovatyi-Kogan, N.R. Ikhsanov  (2014) A new look at anomalous X-ray pulsars. Astron Rep 58:217–227


\bibitem{bk06} G.S. Bisnovatyi-Kogan (2006) Binary and recycled pulsars: 30 years after observational discovery. Phys Usp 49:53–67

\bibitem{bkch79} G.S. Bisnovatyi-Kogan, V.M. Chechetkin (1979) Nonequilibrium shells of neutron stars and their role in sustaining X-ray emission and nucleosythesis. Phys Usp 22:89–108


\bibitem{lx00} D.R. Lorimer, K.M. Xilouris (2000)
PSR J1907+0918: A young radio pulsar near SGR 1900+14 AND G42.8+0.6. Astrophysical Journal, 545:385-389

\bibitem{wk99}  P.M. Woods, C. Kouveliotou, J. van Paradijs, M.H. Finger, C. Thompson, et al (1999) Variable spin-down in the Soft amma Repeater SGR 1900+14 and correlations with burst activity. ApJ 524:L55–L58

\bibitem{xmm06} S. Mereghetti, P. Esposito2, A. Tiengo et al. (2006) The first XMM-Newton observations of the Soft Gamma-ray Repeater SGR 1900+14.  Astrophys J 653:1423-1428

\bibitem{gmt06} D. Götz, S. Mereghetti, A. Tiengo, and P. Esposito (2006) Magnetars as persistent hard X-ray sources: INTEGRAL discovery of a hard tail in SGR 1900+14.
Astron Ap 449, L31–L34 

\bibitem{emt07} P. Esposito, S. Mereghetti, A. Tiengo et al. (2007) Five years of SGR 1900+14 observations with BeppoSAX. Astron Astrophys 461:605-612

\bibitem{irm08} G.L. Israel, P. Romano, V. Mangano, et al. (2008) A SWIFT gaze into the 2006 March 29 burst forest of SGR 1900+14. Astrophysical Journal 685: 1114-1128

\bibitem{tck12}
 S.P. Tendulkar, P.B. Cameron, S.R. Kulkarni (2012)
 Proper motions and origins of SGR1806-20 and SGR1900+14.
 Astrophysical Journal 761:76 (11pp)

\bibitem{tb19} T. Tamba, A. Bamba, H. Odaka,
T. Enoto (2019) Temporal and spectral X-ray properties
of magnetar SGR1900+14 derived from
observations with NuSTAR and XMM-Newton. Publ Astron Soc Japan (2019) 71 (5), 90 (1–13)

\bibitem{ykk15} G. Younes, C. Kouveliotou, and V. M. Kaspi (2015)
XMM-NEWTON observations of SGR 1806–20 over seven years following the 2004 giant flare. Astrophysical Journal, 809:165 (8pp)

\bibitem{ybk17}
G.Younes, M.G. Baring, C. Kouveliotou, et al. (2017)
The sleeping monster: NuSTAR Observations of SGR 1806–20, 11 years after the giant flare. Astrophysical Journal, 851:17 (8pp)

\bibitem{smp14}
 M. Stamatikos, D. Malesani, K.L. Page, T. Sakamoto (2014) GRB 140705A: Swift detection of a short burst. GRB Coordinates Network, 16520

\bibitem{ykj17}
G. Younes, C. Kouveliotou, A. Jaodand, et al. (2017)
X-Ray and radio observations of the magnetar SGR J1935+2154 during
its 2014, 2015, and 2016 outbursts. Astrophysical Journal, 847:85 (15pp)

\bibitem{lgr20}
L. Lin, E. Göğüş, O.J. Roberts, et al. (2020)
Fermi/GBM View of the 2019 and 2020 burst active episodes of SGR J1935+2154. Astrophysical Journal Letters 902:L43 (11pp)

\bibitem{ygk20}
G. Younes, T. Güver, C. Kouveliotou, et al. (2020)
NICER view of the 2020 burst storm and persistent emission of SGR 1935+2154. Astrophysical Journal Letters 904:L21 (16pp) 

\bibitem{gbko20}
E. Göğüş, M.G. Baring, C. Kouveliotou, et al (2020) Persistent emission properties of SGR J1935+2154 during its 2020 active episode. Astrophysical Journal Letters 905:L31 (6pp)

\bibitem{bzi22}
A. Borghese, F.C. Zelati, G.L. Israel (2022)
The first 7 months of the 2020 X-ray outburst of the magnetar
SGR J1935+2154. arXiv:2205.04983v1, 10 May 2022

\end{thebibliography}

\documentclass{article}
\usepackage{srcltx}
\usepackage[T2A]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}

\usepackage[a4paper,top=4cm,bottom=3.5cm,left=2.5cm,right=2.5cm,marginparwidth=1.75cm]{geometry}
\usepackage{graphicx}

\begin{document}

\title{{About the observational check of the mechanism of gamma radiation in  Soft Gamma Repeaters (SGR)}}

 \medskip

    \author{ G.S. Bisnovatyi-Kogan \footnote{{Space research institute RAS, Moscow.}}}

\maketitle

                 {\centerline{\bf{Abstract.}}}
\bigskip

 Soft gamma repeaters (SGR) are identified as single neutron stars (NS) inside the Galaxy, or nearby galaxies, with sporadic transient gamma radiation. A total number of discovered SGR, including relative Anomalous X-ray pulsars (AXP), is few tens of objects. Many of them show periodic radiation, connected with NS rotation, with periods 2-12 s. The slow rotation is accompanied by small rate of loss of rotational energy, which is considerably smaller than the observed sporadic gamma ray luminosity, and is many orders less that the luminosity during giant bursts, observed in 4 SGR. Therefore the energy source is usually connected with annihilation of very strong NS magnetic field. Another model is based on release of a nuclear energy stored in the NS nonequilibrium layer. We suggest here an observational test with could distinguish between these two models.

{\it Keywords}: neutron stars; soft gamma repeaters; magnetic field annihilation; nuclear reactions

 \medskip

\section{Introduction}

 Soft gamma repeaters (SGR),are discovered observationally in the year 1979 \cite{mg79b,mg79c,gm79} as a giant burst with recurrent gamma flared, and soon were connected with the discovered short gamma ray bursts (GRB) \cite{mg82} as a separate class of GRB.
 They  are identified now as single neutron stars (NS) inside the Galaxy, or nearby galaxies, with sporadic transient gamma radiation. A total number of discovered SGR, including relative Anomalous X-ray pulsars (AXP), is few tens of objects \cite{ruf17}. Some of them show periodic radiation, connected with NS rotation, with periods 2-12 s.
It seems now that short GRB include two different types of objects: merging of binary NS at cosmological distances, and another class represented by giant bursts in SGR at neighbouring galaxies at distances, when  recurrent sporadic SGR radiation is too faint for registration at the Earth.
Such events had been discussed in \cite{bk99}, and were discovered
in the galaxies M31 (Andromeda) \cite{ma08}, and a galaxy in the M81 group of galaxies \cite{fp07b}, see also \cite{plp05,hbs05}.

 The slow rotation is accompanied by small rate of loss of rotational energy, which is considerably smaller than the observed sporadic gamma ray luminosity, and is many orders less that the luminosity during giant bursts, observed in 4 SGR
 \cite{mg79c,mc99c,mc05,ma99a}. Therefore the energy source is usually connected with annihilation of very strong NS magnetic field \cite{dt92,dt95}. Another model is based on release of a nuclear energy stored in the NS nonequilibrium layer \cite{bkch74,bkinch75,bk92,bk12}. We suggest here an observational test with could distinguish between these two models.

\section{Periodical pulses}

First discovery of FXP 0520-66 by Konus experiment \cite{mg79b,mg79c} was accompanied by observation of periodic pulsation in this source with the period $P\approx 8$ s. in the afterglow radiation of the giant burst, lasting about 200 s. Similar pulsations had been observed in the afterglows of giant bursts it the objects $SGR 1900+14$ with the period $P\approx 5.16$ s. \cite{mc99c,ks99,mg79a}, and SGR 1806-20 with the period $P\approx 7.47$ s. \cite{mc05,pb05}. The periodicity  in the radiation of this source was found earlier \cite{kd98} in one of the recurrent outbursts, indications to such periodicity had been found in \cite{uf93}, and first discovery of this source in the year 1986 was reported in \cite{lf86a,lf86b}. The source SGR 1806-20 was probably the first SGR, in which periodic pulsations had been discovered in one of the recurrent outbursts, which period was later confirmed in the afterglow of the giant burst 27 December 2004, which was the most powerful from four observed giant bursts. Periodic pulsations had been found in the recurrent outbursts of several other SGR  \cite{ruf17,re10,aec11,gbk20}.

Connection of SGR with highly magnetized slowly rotating NS (magnetars) meets several problems. They are connected with observation of low magnetic field SGR \cite{re10,aec11}. Observations of several, quite normal radio pulsars without traces of recurrent gamma outbursts, having magnetar-like periods and magnetic fields \cite{msk03} are not consistent with this model.
These problems are discussed in details in \cite{bk16,bk17}. It is evident, that magnetic origin of gamma ray outbursts in SGR implies a  different behaviour of pulsations, in comparison with nuclear explosion origin \cite{bkch81,bk12,bki14}.We suggest that this difference should be visible in the behaviour of  pulsation phases, what could be derived from observations.

\section{Phase analysis}

Phase analysis of pulsations in rotating magnetized NS, observed as pulsars (radio, optical, X-ray, gamma-ray) is a very powerful tool, permitting to clarify the properties of the sources. In the binary millisecond pulsars with extremely  precise time stability of pulses the timing analysis of pulses gave the best precision results in favour of confirmation of General Relativity (GR) as a true theory of gravitation (see e.g. \cite{bk06}). This stability is connected with a very stable NS rotation, with practically fixed encoring of the magnetic field to the NS body. Such properties permit to exclude the changes connecting with a motion of magnetic field over NS surface, and interpret the visible period variations in timing data to the pulsar motion in the binary system, and different GR effects.

   In SGR the precision of observation is incomparably worse than in radiopulsars, so only sufficiently strong phase shifting in observations could be noticed and measured.
   NS in SGR is considered as single object, so no  motion in binary, and very small GR effects are not be visible in their timing data. The reason for strong phase shift in the SGR pulsations could be expected due to motion of the the burst location over the NS surface from one burst to another one. Such motion is unlikely to be possible if the transient bursts are connected with magnetic events. Extremely high conductivity of the pulsar matter prevents any magnetic field  changes in static situation, and does not permit changes in the large scale magnetic field configuration formed by electrical currents situated deep inside, it quiet region of the star.
The large scale of the magnetic field should determine the position of the bursting spot in all cases, not permitting for its motion over the star surface.
   The shift of the explosion spot over the surface of the NS is possible in the case when the explosion  mechanism  is not connected directly to the magnetic field large scale structure. Such mechanism of explosion, connected with processes in the NS non-equilibrium layer \cite{bkch74} was considered in \cite{bk12,bki14,bk16,bk17}. Earlier it was suggested as a mechanism for formation of gamma ray bursts (GRB) inside the Galaxy \cite{bkinch75,bkch79}. Discovery of cosmological origin of GRB exclude this mechanism for GRB model, but it remains for application to SRG bursts. A considerable phase shift in the gamma ray pulsar timing, between the observed transient bursts, would be a strong argument in favour of the nuclear explosion SGR burst mechanism.

   \section{Phase shift observations in SGR}

   The pulsed radiation of SGR was first discovered in the famous 5 March 1979 burst \cite{mg79b,mg79c}. Subsequently it was visible very distinctly in the afterglows of two other giant bursts  SGR 1900+14 and SGR 1900+14. Analysis of many recurrent bursts in the last two objects had shown the periodic component in their radiation, with almost the same, slightly growing period.

     \subsection{SGR 1900+14}

The pulsations with a period 5.16 s. have been observed in the afterglow of the giant burst during several hundred of seconds.    \cite{mc99c,ks99}. Search for periodic pulsations had been done in recurrent bursts and persistent radiation of this object in subsequent years in the bands from radio \cite{lx00} to optics and hard X-rays \cite{wk99,xmm06,gmt06,emt07,irm08,tck12,tb19}. The pulsation with a corresponding  period have been observed in all bands, except radio.

   \subsection{SGR 1806-20}

 The pulsed signals with the period 7.47 s. had been discovered  in the year 1998 \cite{uf93,kd98}, 6 years before the giant burst in this SGR. After observations of the giant burst \cite{mc05,pb05}, periodic pulsations in this SGR had been visible in its radiation during subsequent observations in several X-ray bands \cite{ykk15,ybk17}.

   \subsection{SGR J1935+2154}

 SGR J1935+2154 was discovered in 2014 \cite{smp14}, and has a spin period $P \sim 3.25$  s.
 It is one of the most active magnetars, showing
outbursts almost every year \cite{ykj17,lgr20}. Its
reactivation date in 2020 April 27,  was observed by several X-ray and gamma-ray instruments when a burst storm and an increase of the persistent X-ray flux were detected \cite{ygk20,gbko20}.
Subsequent activity of this GRB during 7 months was followed by SMM and Chandra missions \cite{bzi22}.

\section{Conclusions}

The three above considered objects were observed during a long time, when several strong outbursts happen. They looks out as best candidate for making phase analysis, which should show the behaviour of the peak activity spot on the NS surface. If it does not change its position on the NS during all the time, than it is most probably connected with the large scale structure of the magnetic field, and witnesses in favour of the magnetar model. If its position changes after strong outburst that its activity is presumably determined by nuclear explosions, connected with existence of non equilibrium layer in  NS  envelope.

\section*{Acknowledgements}

The author is grateful A.S Pozanenko for discussions, and  N.R. Ikhsanov for useful comments.
This work is partially supported by RFBR grant 20-02-00455.

\end{document}